# Predictive Control and Regret Analysis of Non-Stationary MDP with Look-ahead Information

### v1

@@ Abstract
Policy design in non-stationary Markov Decision Processes (MDPs) is inherently challenging due to the complexities introduced by time-varying system transition and reward, which make it difficult for learners to determine the optimal actions for maximizing cumulative future rewards. Fortunately, in many practical applications, such as energy systems, look-ahead predictions are available, including forecasts for renewable energy generation and demand. In this paper, we leverage these look-ahead predictions and propose an algorithm designed to achieve low regret in non-stationary MDPs by incorporating such predictions. Our theoretical analysis demonstrates that, under certain assumptions, the regret decreases exponentially as the look-ahead window expands. When the system prediction is subject to error, the regret does not explode even if the prediction error grows sub-exponentially as a function of the prediction horizon. We validate our approach through simulations, confirming the efficacy of our algorithm in non-stationary environments.
@@ 1 Introduction
Policy design of non-stationary Markov Decision Processes (MDPs) has always been challenging due to the time-varying system dynamics and rewards, so the learner usually suffers from uncertainties of future rewards and transitions. Fortunately, exogenous predictions are available in many applications. For example, in energy systems, look-ahead information is available in the form of renewable generation forecasts and demand forecasts Amin et al. [2019]. It is intuitive to design an algorithm that controls the energy system by utilizing that information to concentrate energy usage in the time frame with the lowest energy price and lower the overall energy cost. To give another example, smart servers can make predictions of future internet traffic from historical data Katris and Daskalaki [2015]. Given that the server tries to minimize the average waiting time of all tasks, if there is only light traffic, the average waiting time will be most reduced by only using the fastest server. However, if the smart server forecasts that there will be heavy traffic in the future, all servers should work to reduce the length of the queue.
However, although policy adaptation in a time-varying environment has been extensively studied [Auer et al., 2008; Richards et al., 2021; Zhang et al., 2024; Gajane et al., 2018], they do not typically take advantage of exogenous predictions. One branch of work focuses on adaptation and relies on periodic reset of the controller [Auer et al., 2008] or prior knowledge of the environment [Richards et al., 2021; Zhang et al., 2024]. However, prior knowledge of the environment is usually difficult to obtain, and periodically resetting the controller generates a linear regret. Another branch of work focuses on predicting the future based on past data. As past data can not accurately reflect the system in the future, they often need a previously specified (small) variation budget [Merlis, 2024; Lee et al., 2024; Gajane et al., 2018; Padakandla et al., 2020] to achieve sublinear regret. Overall, existing works demonstrate significant challenges in achieving sublinear regret under general non-stationary MDP on a single trajectory without assumptions on sublinear variation budget.
@@
Contribution: We propose an algorithm, Model Predictive Dynamical Programming (MPDP), that utilizes predictions of transitional probability and reward function to minimize the dynamic regret under the setting of a single trajectory. We utilize the look-ahead information and the span semi-norm to show that, under certain assumptions, the regret decays exponentially with the length of the prediction horizon when the prediction is error-less. Specifically, we show that the proposed algorithm achieves a regret of $O(T\gamma^{\lfloor k/J\rfloor}D)$, where $T$ is the time horizon, $k$ is the prediction horizon, and $\gamma<1,J,D$ are constants determined by the properties of the MDP. Even when the system prediction is subject to error, we demonstrate that the regret does not explode if the growth rate of error (as function of prediction horizon) is subexponential. To the best of knowledge, this paper is the first paper that explores the use of exogenous predictions in non-stationary MDP without prior knowledge and proposes an algorithm with decaying regret as the prediction horizon increases and the prediction error descreases.
@@ 1.1 Related Works
Non-stationary reinforcement learning Many works have been done for reinforcement learning (RL) in non-stationary environment, in which a learner tries to maximize accumulated reward during its lifetime. For example, Merlis [2024]; Lee et al. [2024] both use past data to predict future system dynamics via Temporal Difference (TD) learning. However, the standard TD methods are established on stationary MDP, so its applications on non-stationary MDP are often not satisfactory, requiring a limited variation budge on a single trajectory [Li et al., 2019b; Lee et al., 2024; Wei and Luo, 2021; Chandak et al., 2020] or among a sequence of episodes [Merlis, 2024; Feng et al., 2023]. Other works reset the controller at the end of each episode [Gajane et al., 2018; Auer et al., 2008] and assume the ground-truth MDP stays the same within each episode. Moreover, a similar line of work actively detects the switch of MDPs by maintaining an estimate of system dynamics and reset the controller when a switch is detected Alami et al. [2023]; Dahlin et al. [2023]. More recently, a new branch of work uses exogenous predictions, but requires pre-trained optimal policies for each potential MDP the learner encounters, and designs the policy through a mixture of those optimal policies Pourshamsaei and Nobakhti [2024]. Compared with those works, we utilize exogenous predictions for controlling a non-stationary MDP instead of generating those predictions from past data and do not require any prior knowledge about those MDPs.
@@
Model predictive control Traditionally, model predictive control is a branch of control theory that, at each time step, calculates a predictive trajectory for the upcoming $k$ time steps and then implements the first control action from this trajectory [Lin et al., 2021]. Some of these works seek to achieve guarantees such as static regret [Agarwal et al., 2019; Simchowitz and Foster, 2020], dynamic regret [Li et al., 2019a; Yu et al., 2020], or competitive ratio Shi et al. [2020]. From a theoretical perspective, extensive research has been conducted on the asymptotic properties of MPC, including stability and convergence, under broad assumptions about the system dynamics [Diehl et al., 2011; Angeli et al., 2012]. Remarkably, Lin et al. [2021] shows that the regret in a linear-time-varying system decays exponentially with the length of the prediction horizon. However, most of those works assume linear and deterministic system dynamics in continuous space. In this paper, we focus on an MDP setting with finite state spaces and stochastic transition kernels, where the tools and algorithm design are much different.
@@ 2 Problem Formulation
We start by introducing the discrete non-stationary Markov Decision Process characterized by the tuple $(\mathcal{S},\mathcal{A},T,\{P_{t}\}_{t=0}^{T},\{r_{t}\}_{t=0}^{T})$. Here $\mathcal{S},\mathcal{A}$ denote the discrete state space and action space, respectively. The set of functions $\{P_{t}(\cdot|s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}$ is the collection of transition probability measures indexed by the state-action pair $(s,a)$ and time step $t$. The set of function $\{r_{t}(s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}$ is the expected instantaneous reward, where $r_{t}(s,a)$ is the deterministic reward function taking value in $[0,1]$ incurred by taking action $a$ at state $s$ and time step $t$. We denote $T$ as the time horizon. Lastly, we use $\pi_{t}(\cdot|\cdot)$ to determine a decision policy at time $t$ which maps $s_{t}$ to $a_{t}$ and use $\boldsymbol{\pi}$ to denote the collection of $\{\pi_{t}\}_{t=0}^{T}$.
The learner does not have access to the transition probability $\{P_{t}\}_{t}$ and reward functions $\{r_{t}\}_{t}$. Instead, we consider a setting that the learner can predict the future and obtain estimates of future transitional probabilities $\hat{P}$ and rewards $\hat{r}$ for $k$ time steps. More specifically, at any time $t$, the learner has reward and transitional probability estimation $\hat{r}_{t+\ell|t},\hat{P}_{t+\ell|t}$ for any $\ell\in\{1,\dots,k\}$. The error of the predictions are characterized in the below definition.
Definition 1 (prediction error).
@@
for all $s,a,t$, where $\left\lVert\cdot\right\rVert_{\text{TV}}$ stands for the total variation norm.
@@
The learner implements an algorithm $\mathrm{ALG}$ that at time $t$, observes $s_{t}$ and is given the prediction $\hat{r}_{t+\ell|t},\hat{P}_{t+\ell|t}$ for any $\ell\in\{1,\dots,k\}$. It outputs a policy at step $t$, denoted as $\pi_{t}$, based on the available information. Given the algorithm $\mathrm{ALG}$, its value function is defined by
@@
We also define the offline optimal, which is the optimal value had the learner known all the future transition and reward precisely. Formally, the optimal value function is defined by
@@
The learner’s objective is to design a policy $\pi$ that utilizes the dynamics forecast to maximize the cumulative reward and minimize the regret
@@
A server needs to allocate resources for tasks and minimize the average wait time for each job. We consider the setting that there are $n$ servers, each has a service rate $\mu_{i}$, and a single queue with a time-varying arrival rate $\lambda_{t}$. The system’s state space is the length of the queue and whether each server is idle or busy, and its action space is the decision of sending a job in the queue to an available server or not doing anything. The transitional probability is determined as follows. At each time, a job arrives at the queue with probability $\lambda_{t}/(\lambda_{t}+\sum_{i}\mu_{i})$. Then, a dispatcher decides whether to send a job in the queue to one of the idle servers or wait till the next time step. Also, the busy servers will complete the job and become idle with probability $(\sum_{i}\mu_{i})/(\lambda_{t}+\sum_{i}\mu_{i})$. For a detailed explanation of the setup, see Jali et al. [2024]. In this example, the job arrival rate $\lambda_{t}$ varies in time, and having a prediction of the future arrival rate can help the dispatcher in making the job assignment decision. Existing methods for future Internet traffic forecasting [Katris and Daskalaki, 2015] can greatly help with the process and reduce the average wait time.
@@
Renewable energy generation is heavily influenced by environmental factors, and its prediction is also subject to error with different prediction horizons. We consider a discrete-time model for electric vehicle (EV) charging. A list of EV arrives at the charging station from time step 00 to $T$. The $i$-th EV arrives at the station at time $a_{i}$, if the station is already full, it would leave the station without charging. If there is an unused charging stand, it sets a departure time $d_{i}$ and an energy demand $e_{i}$. Moreover, each charging stand has a charging rate capacity $\mu$, and the station has an overall charging capacity $C$. The station needs to charge all electric vehicles that arrive by setting a charging rate $r_{i,t}$ for each vehicle $i$ at each time step $t$. The energy demand of each vehicle has to be satisfied, i.e. $\sum_{t=a_{i}}^{d_{i}}r_{i,t}=e_{i}$. Moreover, the charging rate has to be below the threshold of each charging stand and the station, i.e. for every $t$, $r_{i,t}\in[0,\mu],\sum_{i=1}^{3}r_{i,t}<C$. For a detailed explanation of the setup, see Chen et al. [2022]. Moreover, the energy price fluctuates with energy supply and demand. Since the energy price fluctuates with time $t$, the station wants to minimize the total cost of energy while still fulfilling the energy demand of all vehicles before their departures. In this setup, the state space is the energy demand of each EV, the energy price, and the time frame in which those demands have to be met. The action space is the decision of whether to supply the required amount of energy now or to wait for a later time. As both weather and energy demand can be forecasted [Qu et al., 2013; Amin et al., 2019], learners can utilize this information to adjust their policies.
@@ 3.1 Span Semi-norm and Bellman Operator
The properties of span semi-norm is briefly described in Proposition A.1 in Appendix A. More description can be found in Puterman [1994]. We use span semi-norm to quantify our approximation of the $Q$ functions. Since we use $\arg\max$ on the $Q$ function to determine which action the learner should take, shifting the entire $Q$ function by a constant does not affect the policy. Since span semi-norm has this property (see Proposition A.1 in Appendix A), we use span semi-norm to quantify the deviation of our estimated $Q$ function from the ground-truth $Q$ function.
@@
We are now ready to introduce the definition of $J$-stage contraction.
@@
We say an sequence of operators $\{F_{t}\}_{t\in\{1,\dots,T\}}:\mathcal{S}\rightarrow\mathcal{S}$ is a $J$-stage span contraction if exists $\alpha\in(0,1)$, such that
$\operatorname{sp}(F_{t}\circ\cdots\circ F_{t+J}u-F_{t}\circ\cdots\circ F_{t+J}v)<\alpha\operatorname{sp}(u-v),$
@@
There exists $J\in\mathbb{Z}^{+}$ such that for any pair of policies $\boldsymbol{\pi}_{1}=\{\pi_{h}^{1}\}_{h=0}^{T},\boldsymbol{\pi}_{2}=\{\pi_{h}^{2}\}_{h=0}^{T}$ and time step $t\in\{0,\dots,T-J\}$,
$\begin{split}\eta(\boldsymbol{\pi}_{1},\boldsymbol{\pi}_{2})=&\min_{s_{1},s_{2}\in\mathcal{S}}\sum_{j\in\mathcal{S}}\min\bigg{\{}P_{t}^{\pi_{t}^{1}}\circ\dots\circ P_{t+J}^{\pi_{t+J}^{1}}(j|s_{1}),P_{t}^{\pi_{t}^{2}}\circ\dots\circ P_{t+J}^{\pi_{t+J}^{2}}(j|s_{2})\bigg{\}}>0.\end{split}$
The above assumption is similar to the uniform ergodicity assumption in Yu and Mannor [2009]; Li et al. [2019b]. To see this, we note that Assumption III.1 of Yu and Mannor [2009] also requires ergodicity among states in a non-stationary environment under any pairs of policies. This assumption implies that the effect of any mistake decays exponentially with the number of passing time steps. We point out that, if 1 is not satisfied, it would imply that there exist certain situations within the MDP such that, for any $k$ and $T$, there exists an non-stationary MDP that violates Assumption 1, such that no matter any algorithm would have expected regret of $O(T)$. We give a counter-example of such non-stationary MDP in F.1.
In the next proposition, we introduce how 1 establishes contraction, which we will later use to show the decay of regret with respect to the prediction horizon.
@@
For any non-stationary MDP satisfying Assumption 1, $L_{t}$ defined in (3) is a $J$-stage contraction operator with contraction coefficient
$\gamma=1-\min_{\pi_{1},\pi_{2}}\eta(\pi_{1},\pi_{2}).$
The proof of Proposition 3.1 is deferred to Appendix A.

### v2

@@ Abstract
Policy design in non-stationary Markov Decision Processes (MDPs) is inherently challenging due to the complexities introduced by time-varying system transition and reward, which make it difficult for learners to determine the optimal actions for maximizing cumulative future rewards. Fortunately, in many practical applications, such as energy systems, look-ahead predictions are available, including forecasts for renewable energy generation and demand. In this paper, we leverage these look-ahead predictions and propose an algorithm designed to achieve low regret in non-stationary MDPs by incorporating such predictions. Our theoretical analysis demonstrates that, under certain assumptions, the regret decreases exponentially as the look-ahead window expands. When the system prediction is subject to error, the regret does not explode even if the prediction error grows sub-exponentially as a function of the prediction horizon. We validate our approach through simulations and confirm its efficacy in non-stationary environments.
@@ 1 Introduction
Policy design of non-stationary Markov Decision Processes (MDPs) has always been challenging due to the time-varying system dynamics and rewards, so the learner usually suffers from uncertainties of future rewards and transitions. Fortunately, exogenous predictions are available in many applications. For example, in energy systems, look-ahead information is available in the form of renewable generation forecasts and demand forecasts (Amin19). It is intuitive to design an algorithm that controls the energy system by utilizing that information to concentrate energy usage in the time frame with the lowest energy price and lower the overall energy cost. To give another example, smart servers can make predictions of future internet traffic from historical data (Katris15). Given that the dispatcher tries to minimize the average waiting time of all tasks, if there is only light traffic, the average waiting time will be most reduced by only using the fastest server. If the dispatcher forecasts that there will be heavy traffic in the future, all servers should work to reduce the length of the queue.
However, although policy adaptation in a time-varying environment has been extensively studied (Auer08; Richards21; Ziyi24; Gajane18), they do not typically take advantage of exogenous predictions. One branch of work focuses on adaptation and relies on periodic reset of the controller (Auer08) or prior knowledge of the environment (Richards21; Ziyi24). However, prior knowledge of the environment is usually difficult to obtain, and periodically resetting the controller generates a linear regret. Another branch of work focuses on predicting the future based on past data. As past data can not accurately reflect the system in the future, they often need a previously specified (small) variation budget (merlis20241; Lee24; Gajane18; Padakandla20) to achieve sublinear regret. Overall, existing works demonstrate significant challenges in achieving sublinear regret under general non-stationary MDP on a single trajectory without assumptions on sublinear variation budget.
@@
Contribution: We propose an algorithm, Model Predictive Dynamical Programming (MPDP), that utilizes predictions of transitional probability and reward function to minimize the dynamic regret under the setting of a single trajectory. We utilize the look-ahead information and the span semi-norm to show that, under certain assumptions, the regret decays exponentially with the length of the prediction horizon when the prediction is error-less. Specifically, we show that MPDP achieves a regret of $O(T\gamma^{\lfloor k/J\rfloor}D)$, where $T$ is the time horizon, $k$ is the prediction horizon, $\gamma<1$, and $J,D$ are constants determined by the properties of the MDP. Even when the system prediction is subject to error, we demonstrate that the regret does not explode if the growth rate of error is subexponential as a function of the prediction horizon. To the best of our knowledge, this paper is the first paper that explores the use of exogenous predictions in non-stationary MDP without prior knowledge and proposes an algorithm with decaying regret as the prediction horizon increases and the prediction error descreases.
@@ 1.1 Related Works
Non-stationary reinforcement learning Many works have been done for reinforcement learning (RL) in non-stationary environment, in which a learner tries to maximize accumulated reward during its lifetime. For example, merlis20241; Lee24 both use past data to predict future system dynamics via Temporal Difference (TD) learning. However, the standard TD methods are established on stationary MDP, so its applications on non-stationary MDP require a limited variation budge on a single trajectory (Li19; Lee24; wei21; chandak20; jali2025) or among a sequence of episodes (feng23; moon2024; Zhao2022). For example, jali2025 recently achieves a regret of $\tilde{O}(\sqrt{SA}\Delta_{T}^{\frac{1}{9}}T^{\frac{8}{9}})$, where $S$ is the size of state space, $A$ is the size of action space, $\Delta_{T}$ is the total variation budget, and $T$ is the time horizon. Other works reset the controller at the end of each episode (Gajane18; Auer08) and assume the ground-truth MDP stays the same within each episode. In this line of work, merlis20241 also utilizes prediction in policy design and achieves a regret of $O(H^{3}S^{2}A\sqrt{K}\ln(SATK))$, where $K$ is the number of total episodes. However, this work only allows for changes in either the reward function or the transitional probability while the other stays the same for all episodes, whereas in our setting, we allow for changes in both reward function and transitional probability and achieve an exponentially decaying regret. Another line of work actively detects the switch of MDPs by maintaining an estimate of system dynamics and resetting the controller when a switch is detected (alami23; Dahlin23). More recently, a new branch of work uses exogenous predictions, but requires pre-trained optimal policies for each potential MDP the learner encounters, and designs the policy through a mixture of those optimal policies (Pourshamsaei24). Compared with those works, this paper focuses on a different setting of policy design on a single rollout, which does not allow reset of system at the end of each episode for traditional non-stationary episodic RL. Furthermore, we utilize exogenous predictions for controlling a non-stationary MDP instead of generating those predictions from past data and do not require any prior knowledge about those MDPs.
@@
Model predictive control Traditionally, model predictive control is a branch of control theory that, at each time step, calculates a predictive trajectory for the upcoming $k$ time steps and then implements the first control action from this trajectory (Lin21). Some of these works seek to achieve guarantees such as static regret (Agarwal19; Simchowitz20), dynamic regret (Li191; Yu20), or competitive ratio (Shi20). From a theoretical perspective, extensive research has been conducted on the asymptotic properties of MPC, including stability and convergence, under broad assumptions about the system dynamics (Diehl11; Angeli12). Remarkably, Lin21 shows that regret in a linear-time-varying system decays exponentially with the length of the prediction horizon. However, most of those works assume linear and deterministic system dynamics in continuous space in order to obtain a theoretical regret guarantee. In this paper, we focus on an MDP setting with finite state spaces and stochastic transition kernels, where the tools and algorithm design are very different.
@@ 2 Problem Formulation
We start by introducing the discrete non-stationary Markov Decision Process characterized by the tuple $(\mathcal{S},\mathcal{A},T,\{P_{t}\}_{t=0}^{T},\{r_{t}\}_{t=0}^{T})$. Here $\mathcal{S},\mathcal{A}$ denote the discrete state space and action space, respectively. The set of functions $\{P_{t}(\cdot|s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}$ is the collection of transition probability measures indexed by the state-action pair $(s,a)$ and time step $t$. The set of function $\{r_{t}(s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}$ is the expected instantaneous reward, where $r_{t}(s,a)$ is the deterministic reward function taking value in $[0,1]$ incurred by taking action $a$ at state $s$ and time step $t$. We denote $T$ as the time horizon. Lastly, we use $\pi_{t}(\cdot|\cdot)$ to denote a decision policy at time $t$ that maps $s_{t}$ to $a_{t}$ and use $\bm{\pi}$ to denote the collection of $\{\pi_{t}\}_{t=0}^{T}$.
The learner does not have access to the transition probability $\{P_{t}\}_{t}$ and reward functions $\{r_{t}\}_{t}$. Instead, we consider a setting in which the learner can predict the future and obtain estimates of future transitional probabilities $\hat{P}$ and rewards $\hat{r}$ for $k$ time steps. More specifically, at any time $t$, the learner has reward and transitional probability estimation $\hat{r}_{t+\ell|t},\hat{P}_{t+\ell|t}$ for any $\ell\in\{1,\dots,k\}$. The error of the predictions are characterized in the below definition.
Assumption 1 (prediction error).
@@
for all $s,a,t$, where $\left\lVert\cdot\right\rVert_{\text{TV}}$ denotes the total variation norm.
@@
The learner implements an algorithm $\mathrm{ALG}$ that at time $t$, observes $s_{t}$ and is given the prediction $\hat{r}_{t+\ell|t},\hat{P}_{t+\ell|t}$ for any $\ell\in\{1,\dots,k\}$. It generates a policy at step $t$, denoted as $\pi_{t}$, based on available information. Given the algorithm $\mathrm{ALG}$, its value function is defined by
@@
We also define the offline optimal, which is the optimal value had the learner known all the future transitions and rewards precisely. Formally, the optimal value function is defined by
@@
The learner’s objective is to design a policy $\pi$ that utilizes the dynamics forecast to maximize the cumulative reward and minimize the dynamic regret
@@
A server needs to allocate resources for tasks and minimize the average wait time for each job. We consider the setting that there are $n$ servers, each has a service rate $\mu_{i}$, and a single queue with a time-varying arrival rate $\lambda_{t}$. The system’s state space is the length of the queue and whether each server is idle or busy, and its action space is the decision of sending a job in the queue to an available server or not doing anything. The transitional probability is determined as follows. At each time, a job arrives at the queue with probability $\lambda_{t}/(\lambda_{t}+\sum_{i}\mu_{i})$. Then, a dispatcher decides whether to send a job in the queue to one of the idle servers or wait till the next time step. Also, the busy servers will complete the job and become idle with probability $(\sum_{i}\mu_{i})/(\lambda_{t}+\sum_{i}\mu_{i})$. For a detailed explanation of the setup, see jali24. In this example, the job arrival rate $\lambda_{t}$ varies in time, and having a prediction of the future arrival rate can help the dispatcher in making the job assignment decision. Existing methods for future Internet traffic forecasting (Katris15) can greatly help with the process and reduce the average wait time.
@@
Renewable energy generation is heavily influenced by environmental factors, and its prediction is also subject to error with different prediction horizons. We consider a discrete-time model for electric vehicle (EV) charging. A list of EV arrives at the charging station from time step 0 to $T$. The $i$-th EV arrives at the station at time $a_{i}$, if the station is already full, it would leave the station without charging. If there is an unused charging stand, it sets a departure time $d_{i}$ and an energy demand $e_{i}$. Moreover, each charging stand has a charging rate capacity $\mu$, and the station has an overall charging capacity $C$. The station needs to charge all electric vehicles that arrive by setting a charging rate $r_{i,t}$ for each vehicle $i$ at each time step $t$. The energy demand of each vehicle has to be satisfied, i.e. $\sum_{t=a_{i}}^{d_{i}}r_{i,t}=e_{i}$. Moreover, the charging rate has to be below the threshold of each charging stand and the station, i.e. for every $t$, $r_{i,t}\in[0,\mu],\sum_{i=1}^{3}r_{i,t}<C$. For a detailed explanation of the setup, see Chen22. Moreover, the energy price fluctuates with energy supply and demand. Since the energy price fluctuates with time $t$, the station wants to minimize the total cost of energy while still fulfilling the energy demand of all vehicles before their departures. In this setup, the state space is the energy demand of each EV, the energy price, and the time frame in which those demands have to be met. The action space is the decision of whether to supply the required amount of energy now or to wait for a later time. As both EV parking demands and energy price can be forecasted (jesper20; Catherine22; Kapoor25), learners can utilize this information to adjust their policies.
@@ 3.1 Span Semi-norm and Bellman Operator
The properties of span semi-norm is briefly described in Proposition˜A.1 in Appendix˜A. More description can be found in Puterman94. We use span semi-norm to quantify our approximation of the $Q$ functions. Since we use $\arg\max$ on the $Q$ function to determine which action the learner should take, shifting the entire $Q$ function by a constant does not affect the policy. Since span semi-norm has this property (see Proposition˜A.1 in Appendix˜A), we use span semi-norm to quantify the deviation of our estimated $Q$ function from the ground-truth $Q$ function.
@@
We now introduce the $J$-stage contraction.
@@
We say a sequence of operators $\{F_{t}\}_{t\in\{1,\dots,T\}}:\mathcal{S}\rightarrow\mathcal{S}$ is a $J$-stage span contraction if exists $\gamma\in(0,1)$, such that
$\operatorname{sp}(F_{t}\circ\cdots\circ F_{t+J}u-F_{t}\circ\cdots\circ F_{t+J}v)<\gamma\operatorname{sp}(u-v),$
@@
There exists $J\in\mathbb{Z}^{+}$ such that for the optimal policy $\bm{\pi}^{*}=\{\pi_{h}^{*}\}_{h=0}^{T}$ and any other policy $\bm{\pi}=\{\pi_{h}\}_{h=0}^{T}$ at time step $t\in\{0,\dots,T-J\}$,
$\begin{split}\eta(\bm{\pi},\bm{\pi}^{*})=&\min_{s_{1},s_{2}\in\mathcal{S}}\sum_{j\in\mathcal{S}}\min\bigg\{P_{t}^{\pi_{t}}\circ\dots\circ P_{t+J}^{\pi_{t+J}}(j|s_{1}),P_{t}^{\pi_{t}^{*}}\circ\dots\circ P_{t+J}^{\pi_{t+J}^{*}}(j|s_{2})\bigg\}>0.\end{split}$
The above assumption is similar to the uniform ergodicity assumption in Yu09; Li19. To see this, we note that Assumption III.1 of Yu09 also requires ergodicity among states in a non-stationary environment under any pairs of policies. This assumption implies that the effect of any mistake decays exponentially with the number of passing time steps. In episodic non-stationary RL (e.g., moon2024; Zhao2022), this assumption is automatically satisfied as the environment resets at the end of each episode, so the learner can recover any mistake in future episodes. Further, we point out that if ˜2 is not satisfied, it would imply that there exist certain situations within the MDP such that, for any $k$ and $T$, there exists an non-stationary MDP that violates Assumption 2, such that any algorithm would have expected regret of $O(T)$. We give a counter-example of such non-stationary MDP in ˜F.1.
In the next proposition, we introduce how ˜2 establishes contraction, which we later use to show the decay of regret.
@@
For any non-stationary MDP satisfying Assumption 2, $L_{t}$ defined in equation 3 is a $J$-stage contraction operator with contraction coefficient
$\gamma=1-\min_{\bm{\pi}}\eta(\bm{\pi},\bm{\pi}^{*}).$
The proof of Proposition˜3.1 is deferred to Appendix˜A.